OpenStudy (anonymous):

Please solve!! l2x - (3x - 7)l = 5

OpenStudy (lilg132):

|2x -(3x - 7)| = 5 or |2x -(3x - 7)| = -5 2x - 3x + 7 = 5 2x - 3x = 5 - 7 -x = -2 2 = x or 2x - 3x + 7 = -5 2x - 3x = -5 - 7 2x - 3x = -12 -x = -12 12 = x

OpenStudy (lilg132):

remember to click good answer

OpenStudy (anonymous):

Great! thank you very much, i will be posting more questions so heads up lol xD all points will go to you bro!

OpenStudy (lilg132):

your welcome